Grade 12Algebra

If1/3,a,b,c,d,e,f,g,432are in G.P then find the product of c and e

Dharmendra
9 Years agoGrade 12

1 Answer

Adithi
9 Years ago
in a G.P. , if there are n terms, then the product of kth and (n-k)th terms is always a constant.
in this G.P.,     1/3(432) = ag = bf = ce = d*d
So, c*e = 1/3(432) = 144